ERAM Meeting
-
-
11:00
→
11:45
ERAM statusConveners: Alain Delbart (Université Paris-Saclay (FR)), David Henaff (CEA), Guillaume Eurin (CEA/IRFU/DPhP), Laura-Iuliana Munteanu (CERN)
- 11:00
-
11:05
Mesh pulsing 10mSpeakers: Alain Delbart (Université Paris-Saclay (FR)), David Henaff (CEA), Guillaume Eurin (CEA/IRFU/DPhP), Laura Munteanu (CERN), Laura-Iuliana Munteanu (CERN)
- all mesh pulsing done. ERAMs remaining :
- ERAM-38 : at cabling division. Urgent to have feedback and discuss possibility to remove the connector for which 6 pads are dead and resolder a new one
- ERAM-42 : redo a last mesh pulsing with new cards. If still random 10 rms noise, give it for inspection at cabling division
- ERAM-44 : ready to be glued
- ERAM-45 : 2 dead pads in bottom right corner. Could be glued but wait for possible result of ERAM-38 repair before to decide gluing.
- all mesh pulsing done. ERAMs remaining :
-
11:15
ERAM gluing 5mSpeakers: Alain Delbart (Université Paris-Saclay (FR)), Charlotte Riccio (CEA-Saclay (FR)), David Henaff (CEA), Guillaume Eurin (CEA/IRFU/DPhP), Laura-Iuliana Munteanu (CERN)
- The 2 stiffeners from saclay (sandblasted at Saclay) were used for ERAM-41 and 43
- 3 stiifener at CERN. 2 should be sandblasted again. Is the last one recently found in a good state for gluing ERAM-44 ?
- the O-ring groove was measured by cern metrology. Dimensions are smaller than the ones in the step files. IRFU mechanical division will update the drawings and ask for the resuming of the production of the 6 new stiffeners
- ERAM-44 ready to be glued
-
11:20
ERAM testing 10mSpeakers: Alain Delbart (Université Paris-Saclay (FR)), Charlotte Riccio (CEA-Saclay (FR)), David Henaff (CEA), Guillaume Eurin (CEA/IRFU/DPhP), Laura Munteanu (CERN), Laura-Iuliana Munteanu (CERN)
-
11:30
ERAM AOB 15mSpeakers: Alain Delbart (Université Paris-Saclay (FR)), Charlotte Riccio (CEA-Saclay (FR)), David Henaff (CEA), Guillaume Eurin (CEA/IRFU/DPhP), Laura-Iuliana Munteanu (CERN)
- 16 ERAMs on bot TPC + ERAM-12 (large RC) and 35 (200 kOhm) as spares (at JPARC)
- Global Status of ERAMs production (to be cross checked) for TOP TPC
ERAM-07 : pb seen when bot TPC assembly (not short circuit but too much dust)
ERAM-11
ERAM-14
ERAM-15 : mounting pb on bot TPC --> checked OK at saclay --> usable
ERAM-16 : cc repaired
ERAM-20 : large gain in border
ERAM-27 : THE "surplisingly good" one from bad batch 25-34
ERAM-36 small spreading issues
ERAM-37 small spreading issues
ERAM-39
ERAM-41
ERAM-43 ?
- ERAMs considered "last choices" for TOP TPC
MM1-DLC2 (ERAM-00 in database)
MM1-DLC1
ERAM-25 : bad batch, gain gradient
ERAM-33 : bad batch, gain gradient
ERAM-34 : bad batch, gain gradient
ERAM-38 if not repaired (6 dead pads)
ERAM-42 if not repaired (10 rms noise randomly spread on the global surface)
ERAM-18 : RC/2
- first selection of "first choice" ERAMs for TOP TPC : 12 ERAMs + 2 ? (38 & 42 )
ERAM-07 : pb seen when bot TPC assembly (not short circuit but too much dust)
ERAM-11
ERAM-14
ERAM-15 : mounting pb on bot TPC --> checked OK at saclay --> usable
ERAM-16 : cc repaired
ERAM-20 : large gain in border
ERAM-27 : THE "surplisingly good" one from bad batch 25-34
ERAM-36 small spreading issues
ERAM-37 small spreading issues
ERAM-39
ERAM-41
ERAM-43 ?
- remaining ERAMs = 5 ERAMs
to glue/scan
ERAM-44
ERAM-45
to be produced
ERAM-46
ERAM-47
ERAM-48
- Status of FEC cards available at CERN (LPNHE has 4 FEC more FEC cards to ship to CERN for the last ERAMs)
remaining to be used
15, 23, 28, 31, 59, 60, 73, 75, 76, 83
-
11:00
→
11:45